OpenStudy (anonymous):

Which equation would best help solve the following problem? Holly has a rectangular garden that measures 12 m wide by 14 m long. She wants to increase the area to 255 m2 by increasing the width and length by the same amount. What will be the dimensions of the new garden? A. 255 = (12 – x)(14 – x) B. 255 = (12)(14) C. 255 = (12 + x)(14 + x) D. 255 = (12 + x2)(14)

OpenStudy (anonymous):

Area = 12(14) New area = (12+x)(14+x) 255 = (12+x)(14+x) Expand 255 = 12(14) +12x +14x +x^2 14(12) = 13^2 -1 =168 255= 168 +26x +x^2 Rewrite everything on the left hand side x^2 + 26x +168 - 255 = 0 (168 - 255)=(150 + 18 - 250 -5) Add 150 and -250 (-100 +18 -5) Add 18 and -5 (-100 +13) Add -100 and 13 -87 ____ x^2 +26x -87=0 Use quadratic formula -26 + sqrt(26^2 -4(1)(-87)) all divided by 2 -26 - sqrt(26^2 -4(1)(-87)) all divided by 2 26^2 = 25^2 +25 +26 =625 +50 +1 =676 ______________ 4(87) = 4(90) - 4(3) =360 - 12 = 348 __________________ -26 + sqrt(676+348)all divided by 2 -26 - sqrt(676+348)all divided by 2 676+348 = 650 + 25 + 1 + 350 - 2 =1000 +26 -2 =1024 ____________________________ sqrt(1024) =2^5 = 32 ____________________________ -26 + 32 all divided by 2 -26 - 32 all divided by 2 6 divided by 2 -58 divided by 2 3 <-----(must be positive) -26 _______________ New lengths are (12+3) and (14+3) Shortest dimension (width) is 15. Choice B

OpenStudy (anonymous):

lol wow thank you

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

really @xxraider365

OpenStudy (anonymous):

thank you for telling me @TheRealMeeeee